10 ml of sulphuric acid solution contains 98% by weight of pure acid. Calculate the volume of 2.5 M NaOH solution required to just neutralize the acid.
 

Grade:11

1 Answers

Shreyan Saha
4 years ago
10ml H2SO4 solution = (10×18.4)g solution=(0.98×18.4)g = 18.032 g of pure H2SO4.
18.032 g H2SO4 = 0.184 mol
molarity =(0.184×1000)/10=18.4M
Normality = 18.4×2= 36.8
N1*V1=N2*V2
36.8×10=2.5*x
x=147.2ml